OpenStudy (clamin):

PLEASE HELP!! MEDAL! What affect does adding or subtracting a value "outside" the x have on the graph of the exponential function?

jimthompson5910 (jim_thompson5910):

You mean like going from \(\Large y = 2^x\) to \(\Large y = 2^x+3\) ?

OpenStudy (clamin):

this is the graph

jimthompson5910 (jim_thompson5910):

Did you mark the y intercept of y = 2^x to be 1?

OpenStudy (clamin):

yes

jimthompson5910 (jim_thompson5910):

ok just checking

jimthompson5910 (jim_thompson5910):

the +2 at the end of 2^x means you shift all of 2^x up 2 units so (2^x) + 2 has a y intercept of 3, which is what you have, so it looks good

OpenStudy (clamin):

okay thanks